Find an equation of the line containing the given pair of points.
(1,4) and (3,5)
I found m first 5-4/3-1 = 1/2
Then through other sources told me to use this formula
y-y^1=m(x-x^1)
y-4=1/2(x-1)
y-4=1/2x(1/2)(-1/1)
y=1/2x-1/2+4
y=1/2x-2
But the answer is y=1/2x+7/2
Where did I go wrong? Please could you show me how do this by step by step?
Thank you
JPA

Answer by Alan3354(69443)

Find an equation of the line containing the given pair of points.
(1,4) and (3,5)
I found m first 5-4/3-1 = 1/2
Then through other sources told me to use this formula
y-y^1=m(x-x^1)
y-4=1/2(x-1)
---------------
y-4 = x/2 - 1/2
y = x/2 + 7/2
----------------
y-4=1/2x(1/2)(-1/1) ???
y=1/2x-1/2+4 This is y = (1/2)x - (1/2) + 4 = (1/2)x + 7/2
--------------
A better method, after finding the slope, is to use:
y = mx + b and find b. (Use either point)
4 = (1/2)*1 + b
b = 7/2
--> y = (1/2)x + 7/2
--------------------
The lack of parentheses is confusing, even to you.
y=1/2x-2
But the answer is y=1/2x+7/2